In a certain fraction, the denominator is 4 less than the numerator. If 3 is added to both the numerator and the denominator, the resulting fraction is equal to 5/7. What was the orginal fraction?
Answer by checkley77(12844)

(x+3)/(x-4+3)=5/7
(x+3)/(x-1)=5/7
7(x+3)=5(x-1)
7x+21=5x-5
7x-5x=-5-21
2x=-26
x=-26/2
x=-13
Proof:
(-13+3)/(-13-4+3)=5/7
-10/-14=5/7
5/7=5/7